Jack has just turned 40 years old. He has currently accumulated $55,000 toward his planned retirement at age 60. He wants to accumulate enough money over the next 20 years to provide for a 25-year retirement annuity of $100,000 at the beginning of each year, starting with his 60th birthday. He plans to save $8,000 at the end of each year for the next 10 years. What equal amount must be saved at the end of years 11 through 20 to meet this objective. The interest rate for the first 10 years will be 5%. After that time, the interest rate is expected to be 8%. Please show all steps... I am using the BAII Professional plus...
Answer #1

Step 1:

Set mode to END

PV=55000

PMT=8000

I/Y=5%

N=10

CPT FV=190212.34

Step 2:

Set mode to BGN

N=25

I/Y=8%

PMT=-100000

FV=0

CPT PV=1152875.83

Step 3:

Set mode to END

FV=1152875.83

N=10

I/Y=8%

PV=-190212.34

CPT PMT=51235.18